OpenStudy (anonymous):

I'll give anyone medals and become a fan if they help me. You are flying a kite and want to know its angle of elevation. The string of the kite is 36 meters long and the kite is level with the top of the building you know is 25 meters high. Use an inverse trigonometric function to find the angle of elevation of the kite. Round to the nearest hundredth.

OpenStudy (lucaz):

this situation is a problem with right triangle, the string is the hypothenuse, the height of the building is the opposite side of the angle you want to know

OpenStudy (lucaz):

sin(theta)=opp/hyp sin(theta)=25/36 the inverse function you can use here is the arcsine arcsin(25/36)=theta use a calculator to find the angle theta

OpenStudy (anonymous):

Thanks so much

OpenStudy (lucaz):

no prob